You are considering investing $1,000 in a complete portfolio. The complete portfolio is composed of Treasury bills that pay 5% and a risky portfolio, P, constructed with two risky securities, X and Y. The optimal weights of X and Y in P are 60% and 40%, respectively. X has an expected rate of return of 14%, and Y has an expected rate of return of 10%. The dollar values of your positions in X, Y, and Treasury bills would be _________, __________, and __________, respectively, if you decide to hold a complete portfolio that has an expected return of 8%.
Answer #1

Expected Return of X E(Rx) = 14%

Expected Return Y E(Ry) = 10%

Optimal weight of X (Wx) = 60%

Optimal weight of Y (Wy) = 40%

Thus, Expected return of Risky portfolio would be:

E(Rp) = E(R2) *W, +E(R) *W,

E(R) = 0.14 * .60+ 0.10 * 0.40

E(R) = 0.124 = 12.40%

Risk free rate (Rf) = 5%

Expected Return of Complete portfolio (E(Rc) = 8%

Weight of Risk free assets in complete portfolio = Wf

Weight of Risk portfolio in complete portfolio = Wp = (1-Wf)

Thus,

R *W: +E(R)*(1 – W) = E(R)

0.05 *W: +0.124 * (1 - W.) = 0.08

by solving,

W;= 0.5946

and

W,=1-0.5946 = 0.4054

Thus,

Weight of X in Complete Portfolio = 0.60*0.4054 = 0.2432

Weight of Y in Complete Portfolio = 0.40*4054 = 0.1622

Weight of T-bills in Complete Portfolio = 0.5954

Total Investment in Complete portfolio = $1,000

Thus, In Complete Portfolio,

Dollar value of position in X = 1000*0.2432 = $243.20

Dollar value of position in Y = 1000*0.1622 = $162.20

Dollar value of position in T-bills = 1000*0.5954 = $595.40

Please note: Intermediate calculations rounded off to 4 decimal place. please do comment if have other instructions

Hope this will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.